OpenStudy (anonymous):

Choose the correct simplification of (2xy^2)^2(y^2)^3. Answer 2x^2y^9 2x^2y^10 4x^2y^10 4x^2y^9

OpenStudy (accessdenied):

So, we distribute the exponent in the first: (2xy^2)^2 (y^2)^3. 2^2 * x^2 * (y^2)^2 * (y^2)^3 Then we bring the exponent in since (a^b)^c = a^(bc) 4*x^2 * y^4 * y^6 Then add the exponents since y^a * y^b = y^(a + b) 4x^2 y^10